How do you solve the equation 3 to the power of (x + 3) - 3 to the power of (x) = 78 using the log function
Answer by scott8148(6628)

can't use log function because of subtraction

(3^x)(3^3) - (3^x) = 78

27(3^x) - (3^x) = 78

26(3^x) = 78

3^x = 3

x = 1